OpenStudy (anonymous):

what type of polynomial is x(x-5)

OpenStudy (anonymous):

Please:(

OpenStudy (anonymous):

\(x(x-5)\) is a second order polynomial expression. Expanding gives us \(x(x-5)=x^2-5x\) where \(a=1\), \(b=-5\) and \(c=0\).

OpenStudy (anonymous):

you mean a binomial

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Both. Binomal is a second order polynomial.

OpenStudy (anonymous):

Like bi-cycle